OpenStudy (anonymous):

choose the slope intercept equation on the line that passes through the point (4,-6) and is parallel to y=-1/2x+5

OpenStudy (anonymous):

One second

OpenStudy (anonymous):

Ok here's the work:

OpenStudy (anonymous):

So you have the point (4,-6) and you know that the slope has to be the same in order for the line to be parallel....

OpenStudy (anonymous):

So you set up an equation using 4 for x and -6 for y

OpenStudy (anonymous):

So you have: -6=-1/2(4)+b

OpenStudy (anonymous):

Because you know that -6 which is y is equal to -1/2(the slope) times 4 which is x... but you don't know b(the intercept)

OpenStudy (anonymous):

Next step: -6=-2+b

OpenStudy (anonymous):

Just simplify

OpenStudy (anonymous):

Now add 2 from each side: -4=b

OpenStudy (anonymous):

Now write your equation since you know the slope (-1/2) and the intercept (-4)

OpenStudy (anonymous):

so you get: y=-1/2x-4

OpenStudy (anonymous):

That's your equation... any questions?

OpenStudy (anonymous):

no thank you!!

OpenStudy (anonymous):

No problem!